The data represent the​ time, in​ minutes, spent reading a political blog in a day. Construct a frequency distribution using 5 classes. In the​ table, include the​ midpoints, relative​ frequencies, and cumulative frequencies. Which class has the greatest frequency and which has the least​ frequency? Data Set: 31 12 11 7 5 49 22 38 8 16 2 10 31 37 34 39 37 7 0 41
Answer #1

The data is first arranged in ascending order as:

0, 2, 5, 7, 7, 8, 10, 11, 12, 16, 22, 31, 31, 34, 37, 37, 38, 39, 41, 49

Given the lowest value here is 0 and maximum value is 49, to make 5 classes, we will take the class width as 10 and the required output here is obtained as:

Class Frequency Midpoint Relative Frequency Cumulative Frequency
0 - 9 6 4.5 6/20 = 0.3 6
10 - 19 4 14.5 4/20 = 0.2 6 + 4 = 10
20 - 29 1 24.5 1/20 = 0.05 10 + 1 = 11
30 - 39 7 34.5 7/20 = 0.35 11 + 7 = 18
40 - 49 2 44.5 2/20 = 0.10 18 + 2 = 20

The class with the highest frequency here is 30 - 39 with a frequency of 7 and the class with the lowest frequency here is 20 - 29 with a frequency of 1.